# Building sets, Chow rings, and their Hilbert series

### v1

@@ 4.2. A formula by Aluffi–Marcolli–Nascimento
Our next goal is to give a new and self-contained proof of a recent result by Aluffi, Marcolli, and Nascimento [AMN24, Theorem 1.1]. We follow their notation and write $s(n,k)$ for the (signed) Stirling numbers of the first kind and $S(n,k)$ for the Stirling numbers of the second kind.
@@
The following formula for the Poincaré polynomial of ¯$\overline{\mathcal{M}}_{0,n+1}$ holds:
@@
We start from the formula of Corollary 4.3. Since $P_{\overline{\mathcal{M}}_{0,n+1}}(x)$ is palindromic of degree $n-2$, we obtain
@@
$P_{\overline{\mathcal{M}}_{0,n+1}}(x)=(1-x)^{n}\sum_{k\geq 0}\sum_{\begin{subarray}{c}\lambda\vdash[n+k]\\ \ell(\lambda)=k+1\end{subarray}}x^{n-1+k}\prod_{i=1}^{k+1}\chi_{\lambda_{i}}(x^{-1}).$ (7)
@@
$\sum_{j}x^{j}s(a,a-j)S(a-j,b)=x^{a-b}\sum_{\begin{subarray}{c}\lambda\vdash[a]\\ \ell(\lambda)=b\end{subarray}}\prod_{i=1}^{b}\chi_{\lambda_{i}}(x^{-1}).$ (8)
To see this, notice that both sides count the number of permutations $\sigma\in\mathfrak{S}_{a}$ together with a partition of its set of cycles into exactly $b$ blocks. The parameter $-x$ records the number of cycles of the permutation. Now, applying the identity in Equation (8) to the formula for $P_{\overline{\mathcal{M}}_{0,n+1}}(x)$ obtained in Equation (7) yields the equality
@@ 4.3. Generating functions and recursions
Several different formulas for the above polynomials (and their exponential generating functions) were obtained by Getzler [Get95], Manin [Man95], and Keel [Kee92]. In this section, we show how our Theorem 3.6, when restricted to $\mathsf{M}=\mathsf{K}_{n}$ and $\mathcal{G}=\mathcal{G}_{\min}$, can be used to recover these formulas.
@@
Let $H$ be the exponential generating function of the Poincaré polynomial of ¯$\overline{\mathcal{M}}_{0,n+1}$:
$H=\sum_{n\geq 1}P_{\overline{\mathcal{M}}_{0,n+1}}(x)\frac{t^{n}}{n!}=\sum_{n\geq 1}\mathrm{H}_{\mathsf{K}_{n}}^{\mathcal{G}_{\min}}(x)\frac{t^{n}}{n!}.$
@@
    The incidence algebra equality $-\overline{\chi}^{\mathcal{G}_{\min}}\cdot\mathrm{H}^{\mathcal{G}_{\min}}=\delta$ is equivalent to the compositional formula [Get95, p. 228]
    $H\left(t-\frac{(1+t)^{x}-1-xt}{x(x-1)}\right)=t.$
@@
    The incidence algebra equality $\mathrm{H}^{\mathcal{G}_{\min}}\cdot\overline{\chi}^{\mathcal{G}_{\min}}=-\delta$ is equivalent to the functional equation [Man95, Theorem 0.3.1 (0.7)]
@@
The characteristic polynomial is $\overline{\chi}_{\mathsf{K}_{n}}(x)=(x-2)(x-3)\cdots(x-n+1)$. Therefore, the exponential generating series of $-\overline{\chi}_{\mathsf{K}_{n}}^{\mathcal{G}_{\min}}(x)$ can be rewritten in closed form as
@@
The compositional formula [Sta99, Theorem 5.1.4] can be applied to the identity $-\overline{\chi}^{\mathcal{G}_{\min}}\cdot\mathrm{H}^{\mathcal{G}_{\min}}=\delta$, from which the statement $H(G(t))=t$ follows immediately.
@@
$\mathrm{H}_{\mathsf{K}_{n}}^{\mathcal{G}_{\min}}(x)$ $=\mathrm{H}_{\mathsf{K}_{n-1}}^{\mathcal{G}_{\min}}(x)+x\,\sum_{j=2}^{n-1}\binom{n-1}{j}\mathrm{H}_{\mathsf{K}_{j}}^{\mathcal{G}_{\min}}(x)\cdot\mathrm{H}_{\mathsf{K}_{n-j}}^{\mathcal{G}_{\min}}(x),$ (9)
$=(1+x)\,\mathrm{H}_{\mathsf{K}_{n-1}}^{\mathcal{G}_{\min}}(x)+\frac{x}{2}\,\sum_{j=2}^{n-2}\binom{n}{j}\mathrm{H}_{\mathsf{K}_{j}}^{\mathcal{G}_{\min}}(x)\cdot\mathrm{H}_{\mathsf{K}_{n-j}}^{\mathcal{G}_{\min}}(x),$ (10)
@@
The recursion in Equation (9) (and thus, as we will explain below, also Equation (10)) follows from the functional equations proved in the previous statement: this proof is carried out in [Man95, Corollary 0.3.2]. We take the opportunity, however, to prove it combinatorially via the FY-monomials. Consider an FY-monomial with associated $\mathcal{G}_{\min}$-nested set $S$, and let $G\in S$ be the smallest flat whose nontrivial part contains $n$, if it exists. Let $F_{1},\dots,F_{k}$ be the elements of $S_{<G}$ and $H_{1},\dots,H_{h}$ be the elements of $S_{\not\leq G}$, so that the FY-monomial is $\prod_{i=1}^{k}x_{F_{i}}^{e_{i}}x_{G}^{e}\prod_{j=1}^{h}x_{H_{j}}^{e_{j}}$. Notice that $\prod_{i=1}^{k}x_{F_{i}}^{e_{i}}x_{G\setminus\{n\}}^{e-1}$ is an FY-monomial for the complete induced subgraph on $G\setminus\{n\}$ and $\prod_{j=1}^{h}x_{H_{j}\vee G}^{e_{j}}$ is an FY-monomial for the matroid $\mathsf{K}_{n}/G$ with respect to the minimal building set. The first summand of Equation (9) corresponds to FY-monomials where such $G$ does not exist; the other summands correspond to the connected flats $G$ of rank $j$ whose nontrivial part contains $n$.
The recursion appearing in Equation (10) was proved by Keel in [Kee92, p. 550]. It can be deduced from (9) by first rearranging the term in the sum corresponding to $j=n-1$:
@@
$=(1+x)\mathrm{H}_{\mathsf{K}_{n-1}}^{\mathcal{G}_{\min}}(x)+x\,\sum_{j=2}^{n-2}\binom{n-1}{j}\mathrm{H}_{\mathsf{K}_{j}}^{\mathcal{G}_{\min}}(x)\cdot\mathrm{H}_{\mathsf{K}_{n-j}}^{\mathcal{G}_{\min}}(x).$ (11)
Then, using the change of variables $i=n-j$ on the sum of the right hand side, this is equivalent to
@@
$=(1+x)\mathrm{H}_{\mathsf{K}_{n-1}}^{\mathcal{G}_{\min}}(x)+x\,\sum_{j=2}^{n-2}\binom{n-1}{j-1}\mathrm{H}_{\mathsf{K}_{n-j}}^{\mathcal{G}_{\min}}(x)\cdot\mathrm{H}_{\mathsf{K}_{j}}^{\mathcal{G}_{\min}}(x),$ (12)
@@ 5. Conversion between building sets
Given two building sets $\mathcal{G}$ and $\mathcal{G}^{\prime}$ on $\mathcal{L}(\mathsf{M})$, we now explain how to relate $\mathrm{H}_{\mathsf{M}}^{\mathcal{G}}(x)$ and $\mathrm{H}_{\mathsf{M}}^{\mathcal{G}^{\prime}}(x)$. It will suffice to consider the case when $\mathcal{G}$ and $\mathcal{G}^{\prime}$ differ by one element (see the first part of Proposition 5.2). We thus show the following.
@@
For the proof, we need the following description of $D(\mathsf{M},\mathcal{G})$ as the Chow ring of a fan (see [FY04]). The Bergman fan of $\mathsf{M}$ with respect to $\mathcal{G}$ is the fan $\Sigma_{\mathcal{G}}$ in $\mathbb{R}^{E}$ consisting of the cones $\mathbb{R}_{\geq 0}\{\mathbf{e}_{G}:G\in S\}$, one for each $\mathcal{G}$-nested set $S$. Here, we denoted by $\mathbf{e}_{G}$ the sum $\sum_{i\in G}\mathbf{e}_{i}$ of standard basis vectors. The ring $D(\mathsf{M},\mathcal{G})$ is equal to the Chow ring $A^{\bullet}(\Sigma_{\mathcal{G}})$ of the fan $\Sigma_{\mathcal{G}}$. We now note the following.
@@
[FM05, Proof of Theorem 4.2] Let $\mathcal{G}\supset\mathcal{H}$ be two building sets of $\mathcal{L}(\mathsf{M})$, and let $G$ be a minimal element of $\mathcal{G}\setminus\mathcal{H}$. Then, $\mathcal{G}^{\prime}:=\mathcal{G}\setminus\{G\}$ is a building set. Moreover, the fan $\Sigma_{\mathcal{G}}$ is the stellar subdivision of the fan $\Sigma_{\mathcal{G}^{\prime}}$ at the face corresponding to the nested set $f(\mathcal{G}^{\prime}|_{G})$ of factors of $G$ in $\mathcal{G}^{\prime}$.
@@
This follows from [Kee92, Theorem 1 in Appendix], which is stated for Chow rings of blow-ups of smooth varieties along smooth varieties satisfying a certain surjectivity condition (which holds for toric varieties). In our case, the stellar subdivision corresponds to the blow-up of the toric variety $X_{\Sigma}$ along the torus-invariant subvariety corresponding to the cone $\sigma$. ∎
@@
This follows from a repeated application of [BEPV24, Theorem 1.6], which states that for any $G^{\prime}\in\mathcal{G}^{\prime}$, the closed star of $\Sigma_{\mathcal{G}^{\prime}}$ at the ray $\mathbb{R}_{\geq 0}\{\mathbf{e}_{G^{\prime}}\}$ is isomorphic to the product $\Sigma_{\mathcal{G}^{\prime}|_{G^{\prime}}}\times\Sigma_{\mathcal{G}^{\prime}/G^{\prime}}$. ∎
@@ 6.1. Failure of log-concavity
Let $\mathsf{M}$ be the uniform matroid of rank $9$ on the ground set $E=[8]\sqcup[\overline{10}]$ of 18 elements, where $[8]=\{1,\dotsc,8\}$ and $[\overline{10}]=\{\overline{1},\dotsc,\overline{10}\}$. Let
@@
which one verifies to be a building set on $\mathsf{M}$ by noting that every interval $[\varnothing,F]$ for a proper flat $F\neq E$ is a Boolean lattice.
@@
By induction, one verifies that blowing up a complete flag in a projective space yields a variety whose Betti numbers are the binomial coefficients. In our case, we have that after step (1) the Betti numbers read $(1,8,28,56,70,56,28,8,1)$. Note that $2\cdot 28-(56+8)=-8<0$. Then, for step (2), when one blows up a point, we note that the Betti numbers change by adding $1$ to each of the entries that are not the first or the last. In our case, one may also deduce this geometric fact from Theorem 5.1. In particular, the final Betti numbers read $(1,8+45,28+45,56+45,\dotsc)$, and at the third entry, we find that log-concavity fails since $073^{2}-53\cdot 101=-24<$ 0.
@@
The geometry in the previous example of blowing up a complete flag in $\mathbb{P}^{8}$ and then blowing up 45 additional general points can be realized as a minimal building set on a matroid in the following way. We do this in two steps, corresponding to the steps (1) and (2) in the geometry described in the previous example. We will use notions in matroid theory such as free extensions and principal extensions by flats, which can be found in [Oxl11, Chapter 7].
@@
(2) Now, we extend the matroid $\mathsf{M}_{9}$ (on a ground set of size $19$) to a matroid (of the same rank, $9$) on $19+45\cdot 9=424$ elements, such that the set of connected flats will consist of those from $\mathsf{M}_{9}$ together with an additional $45$ flats of corank $1$. For a matroid $\mathsf{M}$ of rank $r$ on a ground set $E$, denote by $\mathsf{M}+\mathsf{U}_{r-1,r}$ the matroid on ground set $E\sqcup[r]$ constructed as follows. First, perform ($r-1$)-many times the free extension of $\mathsf{M}$ to obtain $\mathsf{M}+\{1,\dotsc,r-1\}$, and then take the principal extension of the resulting matroid by the flat $\{1,\dotsc,r-1\}$. (That is, we extend $\mathsf{M}$ by $r$-many elements as freely as possible with the sole constraint that the $r$-many elements that were added do not form a basis). Note that if a flat $F\subseteq E\sqcup[r]$ of $\mathsf{M}+\mathsf{U}_{r-1,r}$ properly intersects $[r]$, then $F\cap[r]$ consists of coloops of $F$, since by construction no element of $[r]$ is in the closure of any proper flat of $\mathsf{M}$. Thus, the proper connected flats of $\mathsf{M}+\mathsf{U}_{r-1,r}$ are the proper connected flats of $\mathsf{M}$ and $[r]$. The matroid we seek is $\mathsf{M}_{9}+\mathsf{U}_{8,9}+\dotsb+\mathsf{U}_{8,9}$ (where $+\mathsf{U}_{8,9}$ is performed $45$ times).
@@
We provide a building set on a Boolean matroid for which the Chow polynomial has non-log-concave coefficients. That is, the $h$-vector of the corresponding nestohedron need not be log-concave. For this connection to nestohedra, we point to [PRW08].
@@
$\left(\binom{d+1}{0},\binom{d+1}{0}+\binom{d+1}{1},\binom{d+1}{0}+\binom{d+1}{1}+\binom{d+1}{2},\dotsc\right).$

### v2

@@ 4.2. A formula by Aluffi–Marcolli–Nascimento
Our next goal is to give a new and self-contained proof of a recent result by Aluffi, Marcolli, and Nascimento [3, Theorem 1.1]. We follow their notation and write $s(n,k)$ for the (signed) Stirling numbers of the first kind and $S(n,k)$ for the Stirling numbers of the second kind.
@@
The following formula for the Poincaré polynomial of $\overline{\mathcal{M}}_{0,n+1}$ holds:
@@
We start from the formula of Corollary 4.3. Since $P_{\overline{\mathcal{M}}_{0,n+1}}(x)$ is palindromic of degree n2n-2, we obtain
@@
(7) $P_{\overline{\mathcal{M}}_{0,n+1}}(x)=(1-x)^{n}\sum_{k\geq 0}\sum_{\begin{subarray}{c}\lambda\vdash[n+k]\\ \ell(\lambda)=k+1\end{subarray}}x^{n-1+k}\prod_{i=1}^{k+1}\chi_{\lambda_{i}}(x^{-1}).$
@@
(8) $\sum_{j}x^{j}s(a,a-j)S(a-j,b)=x^{a-b}\sum_{\begin{subarray}{c}\lambda\vdash[a]\\ \ell(\lambda)=b\end{subarray}}\prod_{i=1}^{b}\chi_{\lambda_{i}}(x^{-1}).$
To see this, notice that both sides count the number of permutations $\sigma\in\mathfrak{S}_{a}$ together with a partition of its set of cycles into exactly $b$ blocks. The parameter x-x records the number of cycles of the permutation. Now, applying the identity in Equation (8) to the formula for $P_{\overline{\mathcal{M}}_{0,n+1}}(x)$ obtained in Equation (7) yields the equality
@@ 4.3. Generating functions and recursions
Several different formulas for the above polynomials (and their exponential generating functions) were obtained by Getzler [26], Manin [33], and Keel [31]. In this section, we show how our Theorem 3.6, when restricted to $\mathsf{M}=\mathsf{K}_{n}$ and $\mathcal{G}=\mathcal{G}_{\min}$, can be used to recover these formulas.
@@
Let $H$ be the exponential generating function of the Poincaré polynomial of $\overline{\mathcal{M}}_{0,n+1}$:
$H=H(x,t)=\sum_{n\geq 1}P_{\overline{\mathcal{M}}_{0,n+1}}(x)\frac{t^{n}}{n!}=\sum_{n\geq 1}\mathrm{H}_{\mathsf{K}_{n}}^{\mathcal{G}_{\min}}(x)\frac{t^{n}}{n!}.$
@@
    The incidence algebra equality $\overline{\chi}^{\mathcal{G}_{\min}}\cdot\mathrm{H}^{\mathcal{G}_{\min}}=\delta$ is equivalent to the compositional formula [26, p. 228]
    $H\left(t-\frac{(1+t)^{x}-1-xt}{x(x-1)},\ t\right)=t.$
@@
    The incidence algebra equality $\mathrm{H}^{\mathcal{G}_{\min}}\cdot\overline{\chi}^{\mathcal{G}_{\min}}=-\delta$ is equivalent to the functional equation [33, Theorem 0.3.1 (0.7)]
@@
The characteristic polynomial is $\overline{\chi}_{\mathsf{K}_{n}}(x)=(x-2)(x-3)\cdots(x-n+1)$. Therefore, the exponential generating series of $\overline{\chi}_{\mathsf{K}_{n}}^{\mathcal{G}_{\min}}(x)$ can be rewritten in closed form as
@@
The compositional formula [41, Theorem 5.1.4] can be applied to the identity $\overline{\chi}^{\mathcal{G}_{\min}}\cdot\mathrm{H}^{\mathcal{G}_{\min}}=\delta$, from which the statement $H(G(t))=t$ follows immediately.
@@
(9) $\mathrm{H}_{\mathsf{K}_{n}}^{\mathcal{G}_{\min}}(x)$ $=\mathrm{H}_{\mathsf{K}_{n-1}}^{\mathcal{G}_{\min}}(x)+x\,\sum_{j=2}^{n-1}\binom{n-1}{j}\mathrm{H}_{\mathsf{K}_{j}}^{\mathcal{G}_{\min}}(x)\cdot\mathrm{H}_{\mathsf{K}_{n-j}}^{\mathcal{G}_{\min}}(x),$
(10) $=(1+x)\,\mathrm{H}_{\mathsf{K}_{n-1}}^{\mathcal{G}_{\min}}(x)+\frac{x}{2}\,\sum_{j=2}^{n-2}\binom{n}{j}\mathrm{H}_{\mathsf{K}_{j}}^{\mathcal{G}_{\min}}(x)\cdot\mathrm{H}_{\mathsf{K}_{n-j}}^{\mathcal{G}_{\min}}(x),$
@@
The recursion in Equation (9) (and thus, as we will explain below, also Equation (10)) follows from the functional equations proved in the previous statement: this proof is carried out in [33, Corollary 0.3.2]. We take the opportunity, however, to prove it combinatorially via the FY-monomials. Consider an FY-monomial with associated $\mathcal{G}_{\min}$-nested set $S$, and let $G\in S$ be the smallest flat whose nontrivial part contains $n$, if it exists. Let $F_{1},\dots,F_{k}$ be the elements of $S_{<G}$ and $H_{1},\dots,H_{h}$ be the elements of $S_{\not\leq G}$, so that the FY-monomial is $\prod_{i=1}^{k}x_{F_{i}}^{e_{i}}x_{G}^{e}\prod_{j=1}^{h}x_{H_{j}}^{e_{j}}$. Notice that $\prod_{i=1}^{k}x_{F_{i}}^{e_{i}}x_{G\setminus\{n\}}^{e-1}$ is an FY-monomial for the complete induced subgraph on $G\setminus\{n\}$ and $\prod_{j=1}^{h}x_{H_{j}\vee G}^{e_{j}}$ is an FY-monomial for the matroid $\mathsf{K}_{n}/G$ with respect to the minimal building set. The first summand of Equation (9) corresponds to FY-monomials where such $G$ does not exist; the other summands correspond to the connected flats $G$ of rank $j$ whose nontrivial part contains $n$.
The recursion appearing in Equation (10) can be deduced from (9) by first rearranging the term in the sum corresponding to j=n1j=n-1:
@@
(11) $=(1+x)\mathrm{H}_{\mathsf{K}_{n-1}}^{\mathcal{G}_{\min}}(x)+x\,\sum_{j=2}^{n-2}\binom{n-1}{j}\mathrm{H}_{\mathsf{K}_{j}}^{\mathcal{G}_{\min}}(x)\cdot\mathrm{H}_{\mathsf{K}_{n-j}}^{\mathcal{G}_{\min}}(x).$
Then, using the change of variables i=nji=n-j on the sum of the right-hand side, this is equivalent to
@@
(12) $=(1+x)\mathrm{H}_{\mathsf{K}_{n-1}}^{\mathcal{G}_{\min}}(x)+x\,\sum_{j=2}^{n-2}\binom{n-1}{j-1}\mathrm{H}_{\mathsf{K}_{n-j}}^{\mathcal{G}_{\min}}(x)\cdot\mathrm{H}_{\mathsf{K}_{j}}^{\mathcal{G}_{\min}}(x),$
@@ 5. Conversion between building sets
Given two building sets $\mathcal{G}$ and $\mathcal{G}^{\prime}$ on $\mathcal{L}(\mathsf{M})$ with $\mathcal{G}^{\prime}\subseteq\mathcal{G}$, we now explain how to relate $\mathrm{H}_{\mathsf{M}}^{\mathcal{G}}(x)$ and $\mathrm{H}_{\mathsf{M}}^{\mathcal{G}^{\prime}}(x)$. For simplicity, we assume $\mathsf{M}$ to be a matroid in this section. It will suffice to consider the case when $\mathcal{G}$ and $\mathcal{G}^{\prime}$ differ by one element (see the first part of Proposition 5.2). We thus show the following.
@@
For the proof, we need the following description of $D(\mathsf{M},\mathcal{G})$ as the Chow ring of a fan (see [20]). The Bergman fan of $\mathsf{M}$ with respect to $\mathcal{G}$ is the fan $\Sigma_{\mathcal{G}}$ in $\mathbb{R}^{E}$ consisting of the cones $\mathbb{R}_{\geq 0}\{\mathbf{e}_{G}:G\in S\}$, one for each $\mathcal{G}$-nested set $S$. Here, we denoted by $\mathbf{e}_{G}$ the sum $\sum_{i\in G}\mathbf{e}_{i}$ of standard basis vectors. The ring $D(\mathsf{M},\mathcal{G})$ is equal to the Chow ring $A^{\bullet}(\Sigma_{\mathcal{G}})$ of the fan $\Sigma_{\mathcal{G}}$. We now note the following.
@@
[19, Proof of Theorem 4.2] Let $\mathcal{G}\supset\mathcal{H}$ be two building sets of $\mathcal{L}(\mathsf{M})$, and let $G$ be a minimal element of $\mathcal{G}\setminus\mathcal{H}$. Then, $\mathcal{G}^{\prime}:=\mathcal{G}\setminus\{G\}$ is a building set. Moreover, the fan $\Sigma_{\mathcal{G}}$ is the stellar subdivision of the fan $\Sigma_{\mathcal{G}^{\prime}}$ at the face corresponding to the nested set $f(\mathcal{G}^{\prime}|_{G})$ of factors of $G$ in $\mathcal{G}^{\prime}$.
@@
This follows from [31, Theorem 1 in Appendix], which is stated for Chow rings of blow-ups of smooth varieties along smooth varieties satisfying a certain surjectivity condition (which holds for toric varieties). In our case, the stellar subdivision corresponds to the blow-up of the toric variety $X_{\Sigma}$ along the torus-invariant subvariety corresponding to the cone $\sigma$. ∎
@@
This follows from a repeated application of [10, Theorem 1.7], which states that for any $G^{\prime}\in\mathcal{G}^{\prime}$, the closed star of $\Sigma_{\mathcal{G}^{\prime}}$ at the ray $\mathbb{R}_{\geq 0}\{\mathbf{e}_{G^{\prime}}\}$ is isomorphic to the product $\Sigma_{\mathcal{G}^{\prime}|_{G^{\prime}}}\times\Sigma_{\mathcal{G}^{\prime}/G^{\prime}}$. ∎
@@ 6.1. Failure of log-concavity
Let $\mathsf{M}$ be the uniform matroid of rank $9$ on the ground set $E=[8]\sqcup[\overline{10}]$ of 18 elements, where $8]=\{1,\dotsc,8\}$ and $\overline{10}]=\{\overline{1},\dotsc,\overline{10}\}$. Let
@@
which one verifies to be a building set on $\mathsf{M}$ by noting that every interval $\varnothing,F]$ for a proper flat $F\neq E$ is a Boolean lattice.
@@
By induction, one verifies that blowing up a complete flag in a projective space yields a variety whose Betti numbers are the binomial coefficients. In our case, we have that after step (1) the Betti numbers read $(1,8,28,56,70,56,28,8,1)$. Note that $2\cdot 28-(56+8)=-8<0$. Then, for step (2), when one blows up a point, we note that the Betti numbers change by adding $1$ to each of the entries that are not the first or the last. In our case, one may also deduce this geometric fact from Theorem 5.1. In particular, the final Betti numbers read $(1,8+45,28+45,56+45,\dotsc)$, and at the third entry, we find that log-concavity fails since $73^{2}-53\cdot 101=-24<0$.
@@
The geometry in the previous example of blowing up a complete flag in $\mathbb{P}^{8}$ and then blowing up 45 additional general points can be realized as a minimal building set on a matroid in the following way. We do this in two steps, corresponding to the steps (1) and (2) in the geometry described in the previous example. We will use notions in matroid theory such as free extensions and principal extensions by flats, which can be found in [35, Chapter 7].
@@
(2) Now, we extend the matroid $\mathsf{M}_{9}$ (on a ground set of size $19$) to a matroid (of the same rank, $9$) on $19+45\cdot 9=424$ elements, such that the set of connected flats will consist of those from $\mathsf{M}_{9}$ together with an additional $45$ flats of corank $1$. For a matroid $\mathsf{M}$ of rank $r$ on a ground set $E$, denote by $\mathsf{M}+\mathsf{U}_{r-1,r}$ the matroid on ground set $E\sqcup[r]$ constructed as follows. First, perform (r1r-1)-many times the free extension of $\mathsf{M}$ to obtain $\mathsf{M}+\{1,\dotsc,r-1\}$, and then take the principal extension of the resulting matroid by the flat $\{1,\dotsc,r-1\}$. (That is, we extend $\mathsf{M}$ by $r$-many elements as freely as possible with the sole constraint that the $r$-many elements that were added do not form a basis). Note that if a flat $F\subseteq E\sqcup[r]$ of $\mathsf{M}+\mathsf{U}_{r-1,r}$ properly intersects $[r]$, then $F\cap[r]$ consists of coloops of $F$, since by construction no element of $[r]$ is in the closure of any proper flat of $\mathsf{M}$. Thus, the proper connected flats of $\mathsf{M}+\mathsf{U}_{r-1,r}$ are the proper connected flats of $\mathsf{M}$ and $[r]$. The matroid we seek is $\mathsf{M}_{9}+\mathsf{U}_{8,9}+\dotsb+\mathsf{U}_{8,9}$ (where $\mathsf{U}_{8,9}$ is performed $45$ times).
@@
We provide a building set on a Boolean matroid for which the Chow polynomial has non-log-concave coefficients. That is, the $h$-vector of the corresponding nestohedron need not be log-concave. For this connection to nestohedra, we point to [37].
@@
[…]